# ROCKET-1: Master Open-World Interaction with Visual-Temporal Context Prompting

### v1

@@ Abstract
Vision-language models (VLMs) have excelled in multimodal tasks, but adapting them to embodied decision-making in open-world environments presents challenges. A key issue is the difficulty in smoothly connecting individual entities in low-level observations with abstract concepts required for planning. A common approach to address this problem is through the use of hierarchical agents, where VLMs serve as high-level reasoners that break down tasks into executable sub-tasks, typically specified using language and imagined observations. However, language often fails to effectively convey spatial information, while generating future images with sufficient accuracy remains challenging. To address these limitations, we propose visual-temporal context prompting, a novel communication protocol between VLMs and policy models. This protocol leverages object segmentation from both past and present observations to guide policy-environment interactions. Using this approach, we train ROCKET-1, a low-level policy that predicts actions based on concatenated visual observations and segmentation masks, with real-time object tracking provided by SAM-2. Our method unlocks the full potential of VLMs’ visual-language reasoning abilities, enabling them to solve complex creative tasks, especially those heavily reliant on spatial understanding. Experiments in Minecraft demonstrate that our approach allows agents to accomplish previously unattainable tasks, highlighting the effectiveness of visual-temporal context prompting in embodied decision-making. Codes and demos will be available on the project page: https://craftjarvis.github.io/ROCKET-1.
@@ 1 Introduction
Pre-trained foundation vision-language models (VLMs) (Team et al., 2023; Achiam et al., 2023) have shown impressive performance in reasoning, visual question answering, and task planning (Brohan et al., 2023; Driess et al., 2023; Wang et al., 2023b; Cheng et al., 2024), primarily due to training on internet-scale multimodal data. Recently, there has been growing interest in transferring these capabilities to embodied decision-making in open-world environments. Existing approaches can be broadly categorized into (i) end-to-end and (ii) hierarchical approaches. End-to-end approaches, such as RT-2 (Brohan et al., 2023), Octo (Octo Model Team et al., 2024), LEO (Huang et al., 2023), and OpenVLA (Stone et al., 2023), aim to enable VLMs to interact with environments by collecting robot manipulation trajectory data annotated with text. This data is then tokenized to fine-tune VLMs into vision-language-action models (VLAs) in an end-to-end manner, as illustrated in Figure 2(a). However, collecting such annotated trajectory data is resource-intensive and difficult to scale. Moreover, introducing the action modality risks compromising the foundational abilities of VLMs.
@@
Hierarchical agent architectures typically consist of a high-level reasoner and a low-level policy, which can be trained independently. In this architecture, the “communication protocol” between components defines the capability limits of the agent. Alternative approaches (Wang et al., 2023b, a; Driess et al., 2023) leverage VLMs’ reasoning abilities to zero-shot decompose tasks into language-based sub-tasks, with a separate language-conditioned policy executing them in the environment, refer to Figure 2(b). However, language instructions often fail to effectively convey spatial information, limiting the tasks agents can solve. For example, when multiple homonymous objects appear in an observation image, distinguishing a specific one using language alone may require extensive spatial descriptors, increasing data collection complexity and learning difficulty for the language-conditioned policy. To address this issue, approaches like STEVE-1 (Lifshitz et al., 2023), GROOT-1 (Cai et al., 2023b), and MineDreamer (Zhou et al., 2024) propose using a purely vision-based interface to convey task information to the low-level policy. MineDreamer, in particular, uses hindsight relabeling to train an image-conditioned policy (Lifshitz et al., 2023) for interaction, while jointly fine-tuning VLMs and diffusion models to generate goal images that guide the policy, shown in Figure 2(d). Although replacing language with imagined images as the task interface simplifies data collection and policy learning, predicting future observations requires building a world model, which still faces challenges such as hallucinations, temporal inconsistencies, low-quality generation, and limited temporal scope.
@@
In human task execution, such as object grasping, people do not pre-imagine holding an object but maintain focus on the target object while approaching its affordance. When the object is obscured, humans rely on memory to recall its location and connect past and present visual scenes. This use of visual-temporal context enables humans to solve tasks effectively in novel environments. Building on this idea, we propose a novel communication protocol called visual-temporal context prompting, as shown in Figure 2(e). This allows users/reasoners to apply object segmentation to highlight regions of interest in past and current visual observations and convey interaction-type cues via a set of skill primitives. Based on this, we learn ROCKET-1, a low-level policy that uses visual observations and reasoner-provided segmentations as task prompts to predict actions causally. Specifically, a transformer (Dai et al., 2019) models dependencies between observations, essential for representing tasks in partially observable environments. As a bonus feature, ROCKET-1 can enhance its object-tracking capabilities during inference by integrating the state-of-the-art video segmentation model, SAM-2 (Ravi et al., 2024), in a plug-and-play fashion. Additionally, we propose a backward trajectory relabeling method, which efficiently generates segmentation annotations in reverse temporal order using SAM-2, facilitating the creation of training datasets for ROCKET-1. Finally, we develop a hierarchical agent architecture leveraging visual-temporal context prompting, which perfectly inherits the vision-language reasoning capabilities of foundational VLMs. Experiments in Minecraft demonstrate that our pipeline enables agents to complete tasks previously unattainable by other methods, while the hierarchical architecture effectively solves long-horizon tasks.
@@ 3 Methods
To train ROCKET-1, we prepare interaction trajectory data in the format: , where is the visual observation at time , is a binary mask highlighting the object in for future interaction, denotes the interaction type, and is the action. If both and are zeros, no region is highlighted at . As shown in Figure 3, ROCKET-1 is formalized as a conditioned policy, , which takes a sequence of observations and object-segmented interaction regions to causally predict actions. To effectively encode spatial information, inspired by Zhang et al. (2023), we concatenate the observation and object segmentation pixel-wise into a 4-channel image, which is processed by a visual backbone for deep fusion, followed by an attention pooling layer:
We extend the input channels of the first convolution layer in the pre-trained visual backbone from 3 to 4, initializing the new parameters to zero to minimize impact during early training. A TransformerXL (Dai et al., 2019; Baker et al., 2022) module is then used to model temporal dependencies between observations and incorporate interaction type information to predict the next action :
@@
We seek to build a dataset for training ROCKET-1. The collected trajectory data typically contains only observations and actions. To generate object segmentations and interaction types for each frame, we propose a hindsight relabeling technique (Andrychowicz et al., 2017) combined with an object tracking model (Ravi et al., 2024) for efficient data labeling. We first define a set of interactions and identify frames where interaction events occur, detected using a pre-trained vision-language model, such as Achiam et al. (2023). Then, we traverse the trajectory in reverse order, segmenting interacting objects in frame via an open-vocabulary grounding model (Liu et al., 2023). Finally, SAM-2 (Ravi et al., 2024) is used to track and generate segmentations for frames , where is the window length.
For Minecraft, we use contractor data (Baker et al., 2022) from OpenAI, consisting of 1.6 billion frames of human gameplay. This dataset includes meta information for each frame, recording interaction events such as kill entity, mine block, use item, interact, craft, and switch, eliminating the need for vision-language models to detect events. We observed that interacting objects are often centered in the previous frame, allowing the use of a fixed-position bounding box and point with the SAM model for segmentation, replacing open-vocabulary grounding models. We also introduced an additional interaction type, navigate, where significant player displacement over a trajectory identifies the most prominent object in the final frame as the target. As shown in Figure 4, the entire labeling process can be automated using SAM-2.
@@
Completing complex long-horizon tasks in open-world environments requires agents to have strong commonsense knowledge and visual-language reasoning, both of which are strengths of modern VLMs. As shown in Figure 5, we design a hierarchical agent architecture comprising GPT-4o (Achiam et al., 2023), Molmo (Deitke et al., 2024), SAM-2 (Ravi et al., 2024), and ROCKET-1. GPT-4o decomposes tasks into object interactions based on an observation , leveraging its extensive knowledge and reasoning abilities. Since GPT-4o cannot directly output the object masks, Molmo generates coordinates for the described objects. SAM-2 then produces the object mask from these coordinates and efficiently tracks objects in subsequent observations. ROCKET-1 uses the generated masks and interaction types from GPT-4o to engage with the environment. Due to the high computational cost, GPT-4o and Molmo run at lower frequencies, while SAM-2 and ROCKET-1 operate at the environment’s frequency.
| Hyperparameter | Value |
| Input Image Size | |
| Visual Backbone | EfficientNet-B0 (4 channels) |
| Policy Transformer | TransformerXL |
| Number of Policy Blocks | |
| Hidden Dimension | |
| Trajectory Chunk size | |
| Dropout Rate | |
| Optimizer | AdamW |
| Weight Decay | |
| Learning Rate | |
| Warmup Steps |
@@ 4 Results and Analysis
We first provide a detailed overview of the experimental setup, including the benchmarks, baselines, and implementation details. We then explore ROCKET-1 ’s performance on both short-horizon and long-horizon tasks. Finally, we conduct comprehensive ablation studies to validate the rationale behind our design choices.
| Method | Prompt | Hunt | Mine | Interact | Navigate | Tool | Place | Overall | ||||||
| VPT-bc | N/A | 13 | 16 | 0 | 13 | 3 | 31 | 0 | 9 | 0 | 0 | 0 | 0 | 7 |
| STEVE-1 | Human | 0 | 6 | 0 | 69 | 0 | 3 | 0 | 31 | 91 | 6 | 16 | 0 | 19 |
| GROOT-1 | Human | 9 | 22 | 0 | 6 | 3 | 6 | 0 | 3 | 47 | 13 | 3 | 0 | 9 |
[…]
| ROCKET-1 | Human | |||||||||||||
@@ 4.1 Experimental Setup
Briefly, we present ROCKET-1 ’s model architecture, hyperparameters, and optimizer configurations in a Table 1. During training, each complete trajectory is divided into 128-length segments to reduce memory requirements. During inference, ROCKET-1 can access up to 128 frames of past observations. Most training parameters follow the settings from prior works such as Cai et al. (2023b, 2024); Baker et al. (2022).
@@
| Method | Communication Protocol | Policy | |||||||
| DEPS | language | STEVE-1 | |||||||
| MineDreamer∗ | future image | STEVE-1 | - | - | - | ||||
| OmniJarvis | latent code | GROOT-1 | |||||||
[…]
4.2 Performance on Short-Horizon Tasks
@@
To quantitatively assess ROCKET-1 ’s performance on short-horizon tasks, we evaluated it on the Minecraft Interaction Benchmark, with results as illustrated in Table 2. Since ROCKET-1 operates as a low-level policy, it requires a high-level reasoner to provide prompts within a visual-temporal context, driving ROCKET-1 ’s interactions with the environment. We tested two reasoners: (1) A skilled Minecraft human player, who can provide prompts to ROCKET-1 at any interaction moment, serving as an oracle reasoner that demonstrates the upper bound of ROCKET-1 ’s capabilities. (2) A Molmo 72B model (Deitke et al., 2024), where a predefined Molmo prompt is set for each task to periodically select points in the observation as prompts, which are then processed into object segmentations by the SAM-2 model (Ravi et al., 2024). Between Molmo’s invocations, SAM-2’s tracking capabilities offer object segmentations to guide ROCKET-1. For all baselines, humans provide prompts. We found that ROCKET-1 + Molmo consistently outperformed all baselines across all tasks, notably achieving a success rate in the “place oak door on the diamond block”
task, which no baseline could complete. Furthermore, given that this task is unlikely to appear in the training set, our pipeline demonstrates strong zero-shot generalization capabilities.
4.3 Performance on Long-Horizon Tasks
@@
We compared hierarchical agent architectures based on different communication protocols: (1) language-based approaches, exemplified by DEPS (Wang et al., 2023b); (2) future-image-based methods, represented by MineDreamer (Zhou et al., 2024); (3) latent-code-based methods, as in OmniJarvis (Wang et al., 2024a); and (4) our proposed approach based on visual-temporal context, as illustrated in the Figure 5. For MineDreamer, we used the planner provided by DEPS and MineDreamer as the controller to complete the long-horizon experiment. We evaluated these methods on seven crafting or mining tasks, each requiring long-horizon planning: obtaining a wooden pickaxe (3600), furnace (6000), shears (12000), diamond (24000), steak (6000), obsidian (24000), and pink wool (6000), where the numbers in parentheses represent the time limit for each task. In the first five tasks, the agent starts from scratch, while for the obsidian task, we provided an empty bucket and a diamond pickaxe in advance, and for the pink wool task, we provided shears beforehand. Taking the obsidian task as an example, the player must first locate a nearby water source, fill the bucket, find a nearby lava pool, pour the water to form obsidian, and finally switch to the diamond pickaxe to mine the obsidian. We found that our approach significantly improved success rates on the first five tasks, particularly achieving a increase in the steak task. For the last two tasks, all previous baseline methods failed, while our approach achieved a success rate on the wool dyeing task. Screenshots are presented in Figure 7.
| Variants | P-GAP | FPS |
| baseline (w/o sam2) | ||||
| baseline (w/o sam2) | ||||
| +sam2_tiny | ||||
| +sam2_small | ||||
| +sam2_base_plus | ||||
| +sam2_large |
4.4 What Matters for ROCKET-1?

### v3

# ROCKET-1: Mastering Open-World Interaction with Visual-Temporal Context Prompting

@@ Abstract
Vision-language models (VLMs) have excelled in multimodal tasks, but adapting them to embodied decision-making in open-world environments presents challenges. One critical issue is bridging the gap between discrete entities in low-level observations and the abstract concepts required for effective planning. A common solution is building hierarchical agents, where VLMs serve as high-level reasoners that break down tasks into executable sub-tasks, typically specified using language. However, language suffers from the inability to communicate detailed spatial information. We propose visual-temporal context prompting, a novel communication protocol between VLMs and policy models. This protocol leverages object segmentation from past observations to guide policy-environment interactions. Using this approach, we train ROCKET-1, a low-level policy that predicts actions based on concatenated visual observations and segmentation masks, supported by real-time object tracking from SAM-2. Our method unlocks the potential of VLMs, enabling them to tackle complex tasks that demand spatial reasoning. Experiments in Minecraft show that our approach enables agents to achieve previously unattainable tasks, with a absolute improvement in open-world interaction performance. Codes and demos are now available on the project page: https://craftjarvis.github.io/ROCKET-1.
@@ 1 Introduction
Pre-trained foundation vision-language models (VLMs) (Team et al., 2023; Achiam et al., 2023) have shown impressive performance in reasoning, visual question answering, and task planning (Brohan et al., 2023; Driess et al., 2023; Wang et al., 2023b; Cheng et al., 2024), primarily due to training on internet-scale multimodal data. Recently, there has been growing interest in transferring these capabilities to embodied decision-making in open-world environments. Existing approaches can be broadly categorized into (i) end-to-end and (ii) hierarchical approaches. End-to-end approaches, such as RT-2 (Brohan et al., 2023), Octo (Octo Model Team et al., 2024), LEO (Huang et al., 2023), and OpenVLA (Stone et al., 2023), aim to enable VLMs to interact with environments by collecting robot manipulation trajectory data annotated with text. This data is then tokenized to fine-tune VLMs into vision-language-action models (VLAs) in an end-to-end manner, as illustrated in Figure 2(a). However, collecting such annotated trajectory data is difficult to scale. Moreover, introducing the action modality risks compromising the foundational abilities of VLMs.
@@
Hierarchical agent architectures typically consist of a high-level reasoner and a low-level policy, which can be trained independently. In this architecture, the “communication protocol” between components defines the capability limits of the agent. Alternative approaches (Wang et al., 2023b, a; Driess et al., 2023) leverage VLMs’ reasoning abilities to zero-shot decompose tasks into language-based sub-tasks, with a separate language-conditioned policy executing them in the environment, refer to Figure 2(b). However, language instructions often fail to effectively convey spatial information, limiting the tasks agents can solve. For example, when multiple homonymous objects appear in an observation image, distinguishing a specific one using language alone may require extensive spatial descriptors, increasing data collection complexity and learning difficulty for the language-conditioned policy. To address this issue, approaches like STEVE-1 (Lifshitz et al., 2023), GROOT-1 (Cai et al., 2023b), and MineDreamer (Zhou et al., 2024) propose using a purely vision-based interface to convey task information to the low-level policy. MineDreamer, in particular, uses hindsight relabeling to train an image-conditioned policy (Lifshitz et al., 2023) for interaction, while jointly fine-tuning VLMs and diffusion models to generate goal images that guide the policy, shown in Figure 2(d). Although replacing language with imagined images as the task interface simplifies data collection and policy learning, predicting future observations requires building a world model, which still faces challenges such as hallucinations, temporal inconsistencies, and limited temporal scope.
@@
In human task execution, such as object grasping, people do not pre-imagine holding an object but maintain focus on the target object while approaching its affordance. When the object is obscured, humans rely on memory to recall its location and connect past and present visual scenes. This use of visual-temporal context enables humans to solve tasks effectively in novel environments. Building on this idea, we propose a novel communication protocol called visual-temporal context prompting, as shown in Figure 2(e). This allows users/reasoners to apply object segmentation to highlight regions of interest in past visual observations and convey interaction-type cues via a set of skill primitives. Based on this, we learn ROCKET-1, a low-level policy that uses visual observations and reasoner-provided segmentations as task prompts to predict actions causally. Specifically, a transformer (Dai et al., 2019) models dependencies between observations, essential for representing tasks in partially observable environments. As a bonus feature, ROCKET-1 can enhance its object-tracking capabilities during inference by integrating the state-of-the-art video segmentation model, SAM-2 (Ravi et al., 2024), in a plug-and-play fashion. Additionally, we propose a backward trajectory relabeling method, which efficiently generates segmentation annotations in reverse temporal order using SAM-2, facilitating the creation of training datasets for ROCKET-1. Finally, we develop a hierarchical agent architecture leveraging visual-temporal context prompting, which perfectly inherits the vision-language reasoning capabilities of foundational VLMs. Experiments in Minecraft demonstrate that our pipeline enables agents to complete tasks previously unattainable by other methods, while the hierarchical architecture effectively solves long-horizon tasks.
@@ 3 Methods
To train ROCKET-1, we prepare interaction trajectory data in the format: , where is the visual observation at time , is a binary mask highlighting the object in for future interaction, denotes the interaction type, and is the action. If both and are zeros, no region is highlighted at . As shown in Figure 3, ROCKET-1 is formalized as a conditioned policy, , which takes a sequence of observations and object-segmented interaction regions to causally predict actions. To effectively encode spatial information, inspired by Zhang et al. (2023), we concatenate the observation and object segmentation pixel-wise into a 4-channel image, which is processed by a visual backbone for deep fusion, followed by an self-attention pooling layer:
We extend the input channels of the first convolution in the pre-trained visual backbone from 3 to 4, initializing the new parameters to s to minimize the gap in early training. A TransformerXL (Dai et al., 2019; Baker et al., 2022) module is then used to model temporal dependencies between observations and incorporate interaction type information to predict the next action :
@@
We seek to build a dataset for training ROCKET-1. The collected trajectory data typically contains only observations and actions . To generate object segmentations and interaction types for each frame, we propose a novel hindsight relabeling technique (Andrychowicz et al., 2017) combined with an object tracking model (Ravi et al., 2024) for automatic data labeling. We first abstract a set of interactions and identify frames where interaction events occur, detected using a pre-trained vision-language model, such as Achiam et al. (2023). Then, we traverse the trajectory in reverse order, segmenting interacting objects in frame via an open-vocabulary grounding model, such as (Liu et al., 2023). Finally, SAM-2 (Ravi et al., 2024) is used to track and generate segmentations for frames , where is the window length.
For Minecraft, we use contractor data (Baker et al., 2022) from OpenAI, consisting of 1.6 billion frames of human gameplay. This dataset includes meta information for each frame, recording interaction events such as kill entity, mine block, use item, interact, craft, and switch, eliminating the need for vision-language models to detect events. We observed that interacting objects are often centered in the previous frame, allowing the use of a fixed-position bounding box and point with the SAM-2 model for segmentation, replacing open-vocabulary grounding models. We also introduced an additional interaction type, navigate. If a player’s movement exceeds a set threshold over a period, they are considered to be approaching an object. The object they face in the segment’s final frame is marked as the target, with SAM-2 applied in reverse to identify it in earlier frames. As shown in Figure 4, the entire labeling process can be totally automated.
@@
Completing complex long-horizon tasks in open-world environments requires agents to have strong commonsense knowledge and do visual-language reasoning, both of which are strengths of modern VLMs. As shown in Figure 5, we design a novel hierarchical agent architecture comprising GPT-4o (Achiam et al., 2023), Molmo (Deitke et al., 2024), SAM-2 (Ravi et al., 2024), and ROCKET-1. GPT-4o decomposes tasks into object interactions based on an observation , leveraging its extensive knowledge and reasoning abilities. Since GPT-4o cannot directly output the object masks, we use Molmo to generate coordinates for the described objects. SAM-2 then produces the object mask from these coordinates and efficiently tracks objects in subsequent observations. ROCKET-1 uses the generated masks and interaction types from GPT-4o to engage with the environment. Due to the high computational cost, GPT-4o and Molmo run at lower frequencies, while SAM-2 and ROCKET-1 operate at the env’s frequency.
Hyperparameter Value Input Image Size Visual Backbone EfficientNet-B0 (4 channels) Policy Transformer TransformerXL Number of Policy Blocks Hidden Dimension Trajectory Chunk size Dropout Rate Optimizer AdamW Learning Rate
@@ 4 Results and Analysis
First, we provide a detailed overview of the experimental setup, including the benchmarks, baselines, and implementation details. We then explore ROCKET-1 ’s performance on basic open-world interactions and long-horizon tasks. Finally, we conduct comprehensive ablation studies to validate the rationale behind our design choices.
Method
Prompt
Hunt
Mine
Interact
Navigate
Tool
Place
Avg
VPT-bc
N/A
0.13
0.16
0.00
0.13
0.03
0.31
0.00
0.09
0.00
0.00
0.00
0.00
0.07
STEVE-1
Human
0.00
0.06
0.00
0.69
0.00
0.03
0.00
0.31
0.91
0.06
0.16
0.00
0.19
GROOT-1
Human
0.09
0.22
0.00
0.06
0.03
0.06
0.00
0.03
0.47
0.13
0.03
0.00
0.09
ROCKET-1
Molmo
ROCKET-1
Human
@@ 4.1 Experimental Setup
Briefly, we present ROCKET-1 ’s model architecture, hyperparameters, and optimizer configurations in Table 1. During training, each complete trajectory is divided into 128-length segments to reduce memory requirements. During inference, ROCKET-1 can access up to 128 frames of past observations. Most training parameters follow the settings from prior works such as Cai et al. (2023b, 2024b); Baker et al. (2022).
@@
Method
Communication Protocol
Policy
DEPS
language
STEVE-1
MineDreamer∗
future image
STEVE-1
OmniJarvis
latent code
GROOT-1
Ours
visual-temporal context
ROCKET-1
4.2 ROCKET-1 Masters Minecraft Interactions
@@
We evaluated ROCKET-1 on the Minecraft Interaction Benchmark, with results as illustrated in Table 2. Since ROCKET-1 operates as a low-level policy, it requires a high-level reasoner to provide prompts within a visual-temporal context, driving ROCKET-1 ’s interactions with the environment. We tested two reasoners: (1) A skilled Minecraft human player, who can provide prompts to ROCKET-1 at any interaction moment, serving as an oracle reasoner that demonstrates the upper bound of ROCKET-1 ’s capabilities. (2) A Molmo 72B model (Deitke et al., 2024), where a predefined Molmo prompt is set for each task to periodically select points in the observation as prompts, which are then processed into object segmentations by the SAM-2 model (Ravi et al., 2024). Between Molmo’s invocations, SAM-2’s tracking capabilities offer object segmentations to guide ROCKET-1. For all baselines, humans provide prompts. We found that ROCKET-1 + Molmo consistently outperformed all baselines, notably achieving a success rate in the “place oak door on the diamond block” task that no baseline can solve.
4.3 ROCKET-1 Supports Long-Horizon Tasks
@@
We compared hierarchical agent architectures based on different communication protocols: (1) language-based approaches, exemplified by DEPS (Wang et al., 2023b); (2) future-image-based methods, represented by MineDreamer (Zhou et al., 2024); (3) latent-code-based methods, as in OmniJarvis (Wang et al., 2024a); and (4) our proposed approach based on visual-temporal context, as illustrated in the Figure 5. For MineDreamer, we used the planner provided by DEPS and MineDreamer as the controller to complete the long-horizon experiment. We evaluated these methods on seven tasks, each requiring long-horizon planning: obtaining a wooden pickaxe (3.6k), furnace (6k), shears (12k), diamond (24k), steak (6k), obsidian (24k), and pink wool (6k), where the numbers in parentheses represent the time limit. In the first five tasks, the agent starts from scratch, while for the obsidian task, we provide an empty bucket and a diamond pickaxe in advance, and for the pink wool task, we provide shears. Taking the obsidian task as an example, the player must first locate a nearby water source, fill the bucket, find a nearby lava pool, pour the water to form obsidian, and finally switch to the diamond pickaxe to mine the obsidian. Our approach significantly improved success rates on the first five tasks, particularly achieving a increase in the steak task. For the last two tasks, all previous baseline methods failed, while our approach achieved a success rate on the wool dyeing task. Figure 7 presents screenshots.
4.4 What Matters for Learning ROCKET-1?